# Elementary commutator method for the Dirac equation with long-range perturbations

### v1

@@ 2.3 A priori exponential decay estimate
$\begin{split}Q&\leq C_{9}\bigl(f^{-1}\theta^{\prime}+f^{-1-\rho}\bigr)\Theta+C_{9}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{9}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.17)
@@
and for that it suffices to consider the second term of (2.10) since the contribution from $|\chi_{a,b}^{\prime}|\mathrm{e}^{\theta}$ is obviously absorbed into the second term of (2.17). However, the second term of (2.10) is obviously absorbed into the right-hand side of (2.17) due to Lemma 2.6 and the Cauchy–Schwarz inequality. Thus we obtain (2.17).
@@
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\bigl((c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr)\Theta\\ &\phantom{{}={}}{}-C_{10}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{10}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$
and we finally discuss the coefficient of the first term on the right-hand side of (2.25). Now we restrict the ranges of parameters. We first retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary, so that for some $c_{2}>0$
$(c_{1}-q_{1})\geq 2c_{2}\ \ \text{uniformly for }|x|\geq 2^{a_{0}},$
or on $\mathop{\mathrm{supp}}\Theta=\mathop{\mathrm{supp}}\chi_{a,b}$
$(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}$
$\geq\min\bigl\{(c_{2}-C_{10}\theta^{\prime})f^{-1}+(c_{2}-C_{10}f^{-\rho})f^{-1},(c_{2}-C_{10}f^{-1})\theta^{\prime}+c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr\}.$
We further retake $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, so that for some $c_{3}>0$
$c_{2}-C_{10}f^{-\rho}\geq c_{3}\ \ \text{and}\ \ c_{2}-C_{10}f^{-1}\geq c_{3}$
on the same subset. Then we take $\mu\in(0,1]$ small enough that for some $c_{4}>0$
$c_{2}-C_{10}\theta^{\prime}>c_{4}.$
Lastly, since we have for some $c_{5}>0$
$c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\geq c_{5}\mu f^{-1-\delta}-C_{10}f^{-1-\rho},$
letting $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, we are done with the proof. ∎
@@
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion, and let $\mu>0$ and $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.8. Then, taking the expectation of the inequality from Lemma 2.8 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
@@
$\bar{\chi}_{a}^{1/2}f^{-1/2}\mathrm{e}^{\mu f}\phi\in\mathcal{H}.$
This implies $\mathrm{e}^{\mu f}\phi\in L^{2}_{-1/2}\subset\mathcal{B}^{*}_{0}$, and thus we are done. ∎
2.4 A priori super-exponential decay estimate
To prove Proposition 2.2 we next present to the following key lemma.
Lemma 2.10.
Let $\lambda\gtrless m_{\pm}$, respectively, and fix any $\kappa>0$ and $\mu>0$ in the definition (2.6) of $\Theta$. Then there exist $c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\Theta-C\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}-(H-\lambda)\gamma(H-\lambda),$
respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
Proof.
The proof is very similar to that of Lemma 2.8, and we present only key steps, omitting details of the computations. Note that here the parameters are slightly different, and we squeeze positivity in a different manner from Lemma 2.8. We only discuss the upper sign. Fix any $\lambda>m_{+}$ and $\kappa,\mu>0$ as in the assertion. We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}\geq 2\kappa-f^{-1}\geq c_{1}$ (2.20)
uniformly in $\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}$, but later we may retake it larger. All below estimates are uniform in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$.
In this proof we set
$\begin{split}Q&=\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}\\ &\phantom{{}={}}{}+(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.21)
and now let us start to compute the left-hand side of the asserted inequality. Similarly to (2.11) and (2.12), here noting (2.20), we can deduce
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(\theta^{\prime}-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.22)
with $h_{jk}$ defined as (2.13). To compute and bound the third term on the right-hand side of (2.22), we set
$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(\theta^{\prime}-f^{-1}),$
and then, similarly to (2.14), it follows that
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{2}Q.\end{split}$ (2.23)
We proceed similarly to (2.15), so that
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(\theta^{\prime}-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.24)
Thus by (2.22) and (2.24) we obtain
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.25)
We can control $Q$ as follows. In fact, similarly to (2.17), we can prove
$\begin{split}Q&\leq C_{6}f^{-1}\Theta+C_{6}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.26)
and hence by (2.25) and (2.26)
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{2}-C_{7}f^{-1})f\Theta-C_{7}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}(H-\lambda)f\Theta(H-\lambda).\end{split}$
Finally by letting $a_{0}\in\mathbb{N}_{0}$ larger if necessary we are done with the proof. ∎
Proof of Proposition 2.2.
Fix $\phi\in\mathcal{B}^{*}_{0}$, $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa,\mu>0$ as in the assertion, and take $a_{0}\in\mathbb{N}_{0}$ as in Lemma 2.10. Taking the expectation of the inequality from Lemma 2.10 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\begin{split}\bigl\|\Theta^{1/2}\phi\bigr\|^{2}&\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}+C_{\nu}2^{-\nu}\|\chi_{b-1,b+1}\mathrm{e}^{\theta}\phi\|^{2}.\end{split}$ (2.27)
$\bigl\|(\bar{\chi}_{a}\mathrm{e}^{\theta})^{1/2}\phi\bigr\|^{2}$ $\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}$ (2.28)
with $\bar{\chi}_{a}$ being from (2.7). We next let $\nu\to\infty$ in (2.28), and then by Lebesgue’s monotone convergence theorem again
$\bar{\chi}_{a}^{1/2}\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{H}.$
This implies the assertion. ∎
@@ 2.5 Absence of super-exponentially decaying eigenstates
A key lemma for the proof of Proposition 2.3 is the following.
@@
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\kappa^{3}\Theta-C\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}$
@@
Again the proof is very similar to those of Lemmas 2.8 and 2.10, and we present only key steps. Only the upper sign is discussed. Fix any $\lambda>m_{+}$, and let $\mu=0$. We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}=2\kappa-f^{-1}\geq c_{1}\kappa$ (2.29)
@@
In this proof we set
$\begin{split}Q&=\kappa^{2}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}p_{j}\\ &\phantom{{}={}}{}+\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.30)
and then, similarly to (2.11) and (2.12), or (2.22), noting (2.29), we can deduce
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(2\kappa+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(2\kappa-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.31)
with $h_{jk}$ being from (2.13). Set
$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(2\kappa-f^{-1}),$
and then, similarly to (2.14) or (2.23), the third term on the right-hand side of (2.31) is bounded as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\kappa f}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\kappa f}\bigr)-C_{2}Q.\end{split}$
We proceed similarly to (2.15) or (2.24) as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\mathrm{i}\kappa\alpha_{f}\bigr)\Xi\mathrm{e}^{2\kappa f}\bigl(\alpha_{k}p_{k}-\mathrm{i}\kappa\alpha_{f}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{4}\kappa\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{4}\kappa^{2}\Xi\mathrm{e}^{2\kappa f}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\kappa^{3}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(2\kappa-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.32)
Thus by (2.31) and (2.32)
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\kappa^{3}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.33)
Similarly to (2.17) or (2.26), we have
$\begin{split}Q&\leq C_{6}\kappa^{2}f^{-1}\Theta+C_{6}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.34)
@@
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\kappa^{3}(c_{2}-C_{7}\kappa^{-1}f^{-1})f\Theta-C_{7}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}\kappa(H-\lambda)f\Theta(H-\lambda).\end{split}$
Letting $a_{0}\in\mathbb{N}_{0}$, we are done with the proof. ∎
@@
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion. Let $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.11, and take the expectation of the inequality from Lemma 2.11 on the state $\chi_{a-2,b+2}\phi$. Then for any $\kappa\geq 1$ and $b>a\geq a_{0}$
$\kappa\|\chi_{a,b}^{1/2}\mathrm{e}^{\kappa f}\phi\|^{2}\leq C\|\chi_{a-1,a+1}\mathrm{e}^{\kappa f}\phi\|^{2}+C2^{-b}\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2}.$ (2.35)
By the assumption the second term to the right of (2.35) vanishes as $b\to\infty$, and hence by Lebesgue’s monotone convergence theorem
$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa f}\|^{2}$ $\leq C\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2},$

### v2

@@ 2.3 A priori exponential decay estimate
$\begin{split}&-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\alpha_{i}\alpha_{k}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}\mathrm{e}^{\theta/2}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\mathrm{e}^{\theta/2}\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)-C_{6}Q\\ &=-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{6}Q\\ &=-\tfrac{1}{4}(\lambda-q+H-\lambda)\Xi\mathrm{e}^{\theta}(\lambda-q+H-\lambda)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{f}p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{6}Q\\ &\geq-\tfrac{1}{4}(\lambda-q)\Xi\mathrm{e}^{\theta}(\lambda-q)+\tfrac{1}{8}\bigl(\partial_{i}(\partial_{i}f)\widetilde{\Xi}\theta^{\prime}\mathrm{e}^{\theta}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{7}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})|f^{-1}-\theta^{\prime}|\Theta+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}+\tfrac{d-1}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{8}Q.\end{split}$ (2.18)
Hence by (2.12), (2.13), (2.17), (2.18), and (2.10) we obtain
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\bigl(\min\{f^{-1},\theta^{\prime}\}+|f^{-1}-\theta^{\prime}|f^{-1}\theta^{\prime}\bigr)\Theta-C_{9}Q\\ &\geq c_{2}\min\{f^{-1},\theta^{\prime}\}\Theta-C_{9}Q.\end{split}$ (2.19)
Note that we have dropped a positive contribution since it is not necessary here. However, it is essential in the proof of Proposition 2.2.
Next we discuss $Q$. We claim that
$\begin{split}Q&\leq C_{10}\bigl(|\theta^{\prime\prime}|+f^{-1-\rho}\bigr)\Theta+C_{10}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{10}(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.20)
@@
For that it suffices to consider the second term of (2.11) since the contribution from $|\chi_{a,b}^{\prime}|\mathrm{e}^{\theta}$ is obviously absorbed into the second term of (2.20). However, the second term of (2.11) is obviously absorbed into the right-hand side of (2.20) due to Lemma 2.5 and the Cauchy–Schwarz inequality. Thus we obtain (2.20).
@@
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\bigl(\min\{f^{-1},\theta^{\prime}\}-C_{11}|\theta^{\prime\prime}|-C_{11}f^{-1-\rho}\bigr)\Theta\\ &\phantom{{}={}}{}-C_{11}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{11}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.21)
and finally it suffices to look at the coefficient of the first term on the right-hand side of (2.21). However, it is straightforward by restricting the ranges of parameters. We first retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary, and then choose $\mu\in(0,1]$ sufficiently small. Then we can verify that on $\mathop{\mathrm{supp}}\Theta=\mathop{\mathrm{supp}}\chi_{a,b}$
$\min\{f^{-1},\theta^{\prime}\}-C_{11}|\theta^{\prime\prime}|-C_{11}f^{-1-\rho}\geq c_{3}\min\{f^{-1},\theta^{\prime}\}.$
Thus we are done with the proof. ∎
@@
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion, and we follow the strategy outlined in Remarks 2.8. Let $\kappa_{0}\in[0,\infty]$ be from (2.9), let $\mu>0$ and $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.7, and choose $\kappa\geq 0$ as in Remarks 2.8. Then, taking the expectation of the inequality from Lemma 2.7 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
@@
$\bar{\chi}_{a}^{1/2}f^{-1/2}\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{H}.$
This implies $\mathrm{e}^{(\kappa+\mu)f}\phi\in L^{2}_{-1/2}\subset\mathcal{B}^{*}_{0}$, and thus a contradiction. We are done. ∎
@@ 2.4 Absence of super-exponentially decaying eigenstates
To prove Proposition 2.2 we present to the following key lemma.
@@
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\kappa^{2}f^{-1}\Theta-C\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}$
@@
The proof is very similar to that of Lemma 2.7, and we present only key steps, omitting details of the computations. Note that here we have to take particular care of dependence on the parameter $\kappa\geq 1$, and we squeeze positivity from a different term from Lemma 2.7. We only discuss the upper sign. Fix any $\lambda>m_{+}$ and $\mu=0$ as in the assertion. We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\min\{\lambda-q_{0},\lambda-q_{0}-q_{1}\}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}=2\kappa-f^{-1}\geq c_{1}\kappa$ (2.24)
@@
Now let us start to compute the left-hand side of the asserted inequality. In this proof we set
$\begin{split}Q&=\kappa^{2}\bigl(\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}+p_{j}\bigl(\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}p_{j}\\ &\phantom{{}={}}+\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.25)
Then, similarly to (2.12) and (2.13), we can deduce, noting also (2.24),
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(2\kappa+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(2\kappa-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{2\kappa f}f^{d-2}p_{i}h_{il}\alpha_{l}\Xi f^{4-2d}\mathrm{e}^{-2\kappa f}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.26)
where $h_{jk}$ is defined as (2.14), and
$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(2\kappa-f^{-1}).$
As for the third term on the right-hand side of (2.26), we repeat the arguments as in (2.15), (2.17) and (2.18), and then we obtain
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{2\kappa f}f^{d-2}p_{i}h_{il}\alpha_{l}\widetilde{\Xi}f^{4-2d}\mathrm{e}^{-2\kappa f}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{2\kappa f}\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(2\kappa-f^{-1})\Theta+\tfrac{\kappa}{2}\widetilde{\Xi}f^{-1}\mathrm{e}^{2\kappa f}-C_{3}Q.\end{split}$ (2.27)
Hence by (2.26) and (2.27) it follows that
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\kappa^{2}f^{-1}\Theta-C_{4}Q.\end{split}$ (2.28)
Finally, similarly to (2.20), we can control $Q$ by Lemma 2.5 as
$\begin{split}Q&\leq C_{5}\kappa^{2}f^{-1-\rho}\Theta+C_{5}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}{}+C_{5}\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.29)
@@
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{3}-C_{6}f^{-\rho})\kappa^{2}f^{-1}\Theta+C_{6}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}{}+C_{6}\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$
Therefore, letting $a_{0}\in\mathbb{N}_{0}$ larger if necessary, we are done with the proof. ∎
@@
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion. Let $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.9, and taking the expectation of the inequality from Lemma 2.9 on the state $\chi_{a-2,b+2}\phi$. Then we have for any $\kappa\geq 1$ and $b>a\geq a_{0}$
$\|\chi_{a,b}^{1/2}\mathrm{e}^{\kappa f}\phi\|^{2}\leq C\|\chi_{a-1,a+1}\mathrm{e}^{\kappa f}\phi\|^{2}+C2^{-b}\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2}.$ (2.30)
$\bigl\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa f}\phi\bigr\|^{2}\leq C\|\mathrm{e}^{\kappa f}\chi_{a-1,a+1}\phi\|^{2},$